# pMixFed: Efficient Personalized Federated Learning through Adaptive Layer-Wise Mixup

### v1

@@ 6 Conclusions
We introduced pMixFed, a dynamic, layer-wise personalized federated learning approach that uses mixup to integrate the shared global and personalized local models. Our approach features adaptive partitioning between shared and personalized layers, along with a gradual transition for personalization, enabling seamless adaptation for local clients, improved generalization across clients, and reduced risk of catastrophic forgetting. We provided a theoretical analysis of pMixFed to study the properties of its convergence. Our experiments on three datasets demonstrated its superior performance over existing PFL methods. Empirically, pMixFed exhibited faster training times, increased robustness, and better handling of data heterogeneity compared to state-of-the-art PFL models. Future research directions include exploring multi-Modal personalization and adapting pMixFed for working on resource-constrained devices
7 Acknowledgment
This work is based upon the work partly supported by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S. Department of Transportation National University Transportation Center) headquartered at Clemson University, Clemson, South Carolina, USA. Any opinions, findings, conclusions, and recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of TraCR, and the U.S. Government assumes no liability for the contents or use thereof.
@@ 8.1 Dynamic Mixup Degree
Among the various partitioning strategies for partial PFL discussed in the main paper and in [34], one of the most widely adopted techniques is assigning higher layers of local model $L_{l,k}^{t}$ to personalization while allowing the base layers $L_{g,k}^{t}$ to be shared across clients as the global model [33, 45, 4]. This design aligns with insights from the Model-Agnostic Meta-Learning (MAML) algorithm [10], which demonstrates that lower layers generally retain task-agnostic, generalized features, while the higher layers capture task-specific, personalized characteristics. Accordingly, in this work, we designate the head of the model as containing personalized information, while the base layers represent generalized information shared across clients  [^4] It should be noted that our method is fully adaptable to different partial PFL designs as well discussed in [34]..
Broadcasting: To achieve a nuanced and gradual transition in the mixing process between the global model and the local model, we define the mix degree $\lambda_{i}$ as follows. The local model’s head $L_{n}$ remains frozen, and the head of the global model is excluded from being shared with the local model. The mix degree for each layer $\lambda_{i}$ increases incrementally based on the mix factor $\mu$, such that as we move toward the base layers, the personalization impact decreases. This dynamic behavior is represented as:
@@
Aggregation: The aggregation stage focuses on preserving the generalized information from the history of previous global models to mitigate the catastrophic forgetting problem. In contrast to the broadcasting process, the primary goal here is to retain the generalized information of the previous global model, which encapsulates the history of all prior models $H\big{|}_{1}^{t-1}{G}$. Therefore, the base layers should predominantly be shared from the global model, particularly during the first round of training, where the updated local models are still underdeveloped. Nonetheless, as we transition to the head, it’s less prominent to transfer knowledge from the previous global models. Figure 8 illustrates how the mixup degree transitions from the head to the base layers.
@@ 8.2 Adaptive Mix-Factor
To enable online, adaptive updates of the mix degree $\lambda_{i}$, we implemented several algorithms, as detailed in Section 5.4 Ablation Study. We observed a clear relationship between the mixup degree and the similar behavior to lr, which is further discussed in Section 9.4.1. Inspired by learning rate schedulers, we introduced the mix-factor $\mu$ to adaptively update the layer-wise mixup degree based on the current communication round $\delta=t/T$ and the relative performance of the current local model $L_{k}^{t}$ compared to the global model $G^{t}$. The Sigmoid function in Figure 8 illustrates how $\mu$ evolves with respect to $\delta$ and accuracy ($Acc$). The best results were achieved when setting $b=1$ and using the square of $Acc$ as an exponent. The rationale behind this approach is that a more experienced, better-performing model should share more information. Specifically, if the local model accuracy $Acc_{l}$ significantly exceeds that of the global model ($Acc_{G}$) in the current round ($t$) such that $Acc\gg 1$, less information is shared from the global model. Conversely, when $10<Acc<$ 1, the global model dominates the parameter updates in both the global and local models. Figure 7 shows the distribution of the calculated $\mu$ across different communication rounds for client 0 in the broadcasting stage. In the broadcasting stage, higher $Acc_{l}$ values result in freezing more layers for personalization, leading to a decrease in $\mu$. Conversely, during the aggregation phase, if the global model accuracy $Acc_{G}$ outperforms the updated local model accuracy $Acc_{l^{\prime}}^{(t+1)}$, more base layers are shared. This is especially important during the first round of training, where local models are less stable. As demonstrated in Section 5.3, our proposed method adapts dynamically to performance drops and high-distribution complexity, adjusting the mixup degree as needed. This adaptability is effective even when applied partially to partial PFL methods such as FedAlt, and FedSim.
@@ 8.3 Model Heterogeneity
9 Theoretical Analysis
In this section, we provide the convergence analysis of pMixFed. Moreover, We compare the aggregation process and the server global model update of the FedSGD algorithm with our proposed mixed aggregation stage in pMixFed. We begin by introducing the key notations and assumptions used throughout the convergence analysis.
Notations:
  • $t\in\{0,\ldots,T-1\}$: communication round index.
  • $\eta_{l}$: learning rate for local update, $\eta_{g}$: learning rate for global updates.
  • $\lambda_{k,i}$: mixup coefficient for client $k$ at layer $i$ in round $t$, $\lambda_{k}$: mixup coefficient for client $k$ in round $t$ (assuming uniform across layers).
  • $G^{(t)}$: global model parameters at round $t$, $L_{k}^{(t)}$: local model parameters of client $k$ at round $t$.
  • $|\mathcal{D}_{k}|$: size of the dataset at client $k$, $|\mathcal{D}|$: total size of datasets across all clients.
  • $\nabla L_{k}^{(t+1)}$: gradient of the local model at client $k$ in round $t$.
We make the following assumptions to establish the convergence properties of pMixFed:
Assumption 1.
Unbiased Gradient Estimation with Bounded Variance. The gradient estimate for local model updates is unbiased and has a bounded variance, i.e.,
$\mathbb{E}[\nabla\mathcal{L}_{k}^{(t)}]=\nabla\mathcal{L}_{k}^{(t)},\quad\text{Var}(\nabla\mathcal{L}_{k}^{(t)})\leq\sigma^{2}.$ (9)
Assumption 2 (Smoothness of Local Objectives).
The local objective functions $L_{k}(\cdot)$ are $L$-smooth, i.e.,
$\small\|\nabla L_{k}^{(t+1)}-\nabla L_{k}^{(t)}\|\leq L\|L_{k}^{(t+1)}-L_{k}^{(t)}\|.$ (10)
Assumption 3 (Bounded Gradients).
The gradients at each client are bounded, i.e.,
$\small\|\nabla L_{k}^{(t+1)}\|\leq G,\quad\forall k,t.$ (11)
These assumptions are standard in convergence analysis and ensure that the optimization process is well-behaved.
9.1 Local Training Convergence
We first analyze the local training updates that occur between communication rounds.
Lemma 1.
Local Model Training Progress. Under Assumptions 2 and 1, after $r$ local updates, the expected loss for client $k$ satisfies:
$\mathbb{E}[\mathcal{L}_{k}^{(t+r)}]\leq\mathcal{L}_{k}^{(t)}+\left(\frac{L_{1}\eta_{l}^{2}}{2}-\eta_{l}\right)\sum_{j=0}^{r-1}\|\nabla\mathcal{L}_{k}^{(t+j)}\|^{2}+\frac{L_{1}r\eta_{l}^{2}\sigma^{2}}{2}.$ (12)
This lemma provides a bound on how the local training process improves the loss function. The bound depends on the learning rate $\eta_{l}$, the smoothness constant $L_{1}$, and the variance of the gradient estimates $\sigma^{2}$.
@@ 9.1 Local Training Convergence
9.2 Global Model Aggregation
Next, we consider the effect of aggregating local models at the server after each communication round.
Lemma 2.
Global Model Aggregation Dynamics. After aggregating the local models, the global model loss changes as follows:
$\mathbb{E}[\mathcal{L}^{(t+1)}]\leq\mathbb{E}[\mathcal{L}^{(t)}]+\eta_{g}\delta^{2},$ (13)
where $\delta^{2}$ bounds the difference between the global and local model parameters before and after aggregation.
This lemma shows that, during aggregation, the global loss increases by a term proportional to the global learning rate $\eta_{g}$ and the parameter variation $\delta^{2}$.
9.3 Convergence of pMixFed
With the above lemmas, we can now derive the convergence properties of pMixFed.
Theorem 1.
Convergence Rate of pMixFed. Under the previously stated assumptions, the expected global loss after a complete round of communication and local training satisfies:
@@ 9.3 Convergence of pMixFed
$\small\mathbb{E}[\mathcal{L}^{(t+1)}]\leq\mathcal{L}^{(t)}+\left(\frac{L_{1}\eta_{l}^{2}}{2}-\eta_{l}\right)\sum_{k=1}^{K}\sum_{j=0}^{r-1}\|\nabla\mathcal{L}_{k}^{(t+j)}\|^{2}+\frac{L_{1}r\eta_{l}^{2}\sigma^{2}}{2}+\eta_{g}\delta^{2}.$ (14)
This theorem demonstrates that the global loss decreases with each communication round, with the convergence depending on the local and global learning rates, the gradient norms, and the parameter variation.
Theorem 2.
Non-Convex Convergence Rate. For non-convex loss functions, pMixFed achieves convergence at the following rate:
$\frac{1}{T}\sum_{t=0}^{T-1}\sum_{k=1}^{K}\|\nabla\mathcal{L}_{k}^{(t)}\|^{2}\leq\mathcal{O}(1/T),$ (15)
where $T$ is the total number of communication rounds.
This final theorem indicates that pMixFed achieves non-convex convergence with a rate of $\mathcal{O}(1/T)$, demonstrating that the algorithm improves over time as the number of communication rounds increases.
9.4 The effect of Aggregating model parameters and gradients on catastrophic forgetting
In FedSGD, the gradients are aggregated and the server will be update the global model according to the aggregated gradients. the FedSGD is sometimes preferred over FedAvg due to its potentially faster convergence. However, it lacks robustness in heterogeneous environments. pMixFed leverages the faster convergence characteristics of FedSGD by incorporating early stopping mechanisms, facilitated by the use of mixup. As demonstrated in Section 5, the mixup factor $\lambda$ functions analogously to an SGD update at the server, even though pMixFed aggregates model weights rather than gradients, similar to FedAvg. Section 9.4 provides a more detailed explanation of this mechanism.
the pMixFed algorithm combines the advantages of FedSGD and FedAvg by aggregating model weights rather than gradients, while still ensuring convergence even in heterogeneous data settings. The incorporation of the mixup mechanism enhances stability, providing faster convergence rates compared to FedSGD, particularly in non-convex settings. Considering local SGD steps $r$ and $\eta_{l}$ as the local learning rate we can further expand the above formulation:
$\begin{split}&G^{t+1}=G^{t}-\eta_{g}\sum_{i=1}^{k}\Omega_{k}\left[L_{k}^{(t+r)}-L_{k}^{(t)}\right]\\ &G^{t+1}=G^{t}-\eta_{g}.\eta_{l}\sum_{i=1}^{k}\Omega_{k}\sum_{j=0}^{r-1}\nabla{L_{k}^{(t+j)}}\end{split}$ (16)
Since in FedSGD, the global model $G^{(t)}$ is fully shared with each local model $L_{k}^{(t)}$ in each communication round t, and $\sum_{i=1}^{k}\Omega_{k}=1$, we can update the problem as a constrained optimization problem.
$G^{t+1}=G^{t}-\eta_{g}.\eta_{l}\ \sum_{j=1}^{r-1}\nabla{G^{(t+j)}}$
$\sum_{j=2}^{r-1}\nabla{G^{(t+j)}=(1-\eta_{g}.\eta_{l})G^{t}-(1+\eta_{g}.\eta_{l})G^{t+1}\ }$
The right term of the equation is the history of global model gradients which could be written as $H\big{|}_{2}^{r-1}\nabla{G^{t}}=a.G^{t}-b.G^{t+1}$. The problem of the FedSGD algorithm is that the gradients are too small to keep the state of the previous global model, leading to catastrophic forgetting in partial models that use gradients in their aggregation stage. We hypothesize that the mixup coefficient $\lambda$ acts similarly to the learning rate $\eta$, suggesting that $\lambda$ plays a role analogous to $\eta$ in FedSGD.
FedSGD Update Rule : In FedSGD, the global model $G^{(t)}$ is updated by aggregating the gradients from all clients:
$\small G^{(t+1)}=G^{(t)}-\eta_{g}\sum_{k=1}^{K}\Omega_{k}\nabla L_{k}^{(t+1)},$ (17)
where $\Omega_{k}=\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}$ is the weight associated with client $k$. Since $L_{k}^{(t+1)}=L_{k}^{(t)}-\eta_{l}\nabla L_{k}^{(t+1)}$, we can rewrite the update as:
$\small G^{(t+1)}=G^{(t)}-\eta_{g}\sum_{k=1}^{K}\Omega_{k}\left(\frac{L_{k}^{(t)}-L_{k}^{(t+1)}}{\eta_{l}}\right).$ (18)
pMixFed Update Rule: In pMixFed, the global model update incorporates mixup coefficients:
$\small G^{(t+1)}=\sum_{k=1}^{K}\Omega_{k}\left[(1-\lambda_{k})L_{k}^{(t+1)}+\lambda_{k}G^{(t)}\right].$ (19)
Assuming $\lambda_{k,i}=\lambda_{k}$ for all layers $i$, and considering that $G^{(t)}$ is sent to all clients at round $t$, we simplify the update to:
$\small G^{(t+1)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}+\lambda_{k}G^{(t)}.$ (20)
9.4.1 Analytical analysis of the Effect of learning rate and mixup degree
To establish the relationship between $\lambda$ and $\eta_{g}$, we align the FedSGD and pMixFed update equations. From Equation (20), rearranged:
$\small G^{(t+1)}-\lambda_{k}G^{(t)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (21)
From Equation (18), rearranged:
$\small G^{(t+1)}=G^{(t)}-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\left(L_{k}^{(t)}-L_{k}^{(t+1)}\right).$ (22)
Assuming $L_{k}^{(t)}$ is replaced with $G^{(t)}$ at round $t$ for FedSGD, we have $L_{k}^{(t)}=G^{(t)}$. Substituting this into Equation (22):
$G^{(t+1)}=G^{(t)}-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\left(G^{(t)}-L_{k}^{(t+1)}\right).$ (23)
Simplifying and Comparing with Equation (20), we see that if: $\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}$. then the updates are analogous.
$G^{(t+1)}=G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (24)
Theorem 3.
Under Assumptions 2 and 3, the mixup coefficient $\lambda_{k}$ in pMixFed acts similarly to the learning rate ratio $\frac{\eta_{g}}{\eta_{l}}$ in FedSGD, such that:$\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}$. This implies that the mixup mechanism in pMixFed can be interpreted as a form of learning rate control analogous to FedSGD.
Proof.
Starting from the pMixFed update in Equation (21):
$G^{(t+1)}-\lambda_{k}G^{(t)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (25)
From the rearranged FedSGD update in Equation (24):
$G^{(t+1)}=G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (26)
Setting the two expressions equal:
$G^{(t)}\left(1-\lambda_{k}\right)+(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}=$
$G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (27)
This simplifies to:
$\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}.$ (28)
Given that $\lambda_{k}$ must lie in the range $[0,1]$, this relationship holds when $\eta_{g}\leq\eta_{l}$. Since both $\eta_{g}$ and $\eta_{l}$ are also constrained within the interval $[0,1]$, our findings are consistent with previous studies [33, 10], which recommend keeping the head unfrozen (i.e., $\eta_{l}\neq 0$).
Remark2. practice, this relationship suggests that by tuning $\lambda_{k}$, we can control the influence of the global model versus the local models in the aggregation process, similar to adjusting the learning rate in FedSGD. This is particularly beneficial in heterogeneous environments where clients may have varying data distributions.
Our theoretical analysis indicates that the mixup coefficient $\lambda_{k}$ in pMixFed plays a role analogous to the learning rate in FedSGD. This equivalence provides a deeper understanding of how pMixFed leverages the strengths of FedSGD while mitigating its weaknesses in heterogeneous settings. By appropriately choosing $\lambda_{k}$, pMixFed can achieve faster convergence and improved robustness.

### v2

@@ 7 Conclusions
We introduced pMixFed, a dynamic, layer-wise personalized federated learning approach that uses mixup to integrate the shared global and personalized local models. Our approach features adaptive partitioning between shared and personalized layers, along with a gradual transition for personalization, enabling seamless adaptation for local clients, improved generalization across clients, and reduced risk of catastrophic forgetting. It should be noted that while Mixup has previously been used in FL (e.g., XORMixup, FEDMIX, FedMix), its role was limited to data augmentation. Similarly, model partitioning has been explored in works like FedAlt and FedSim, but only with fixed cut-off layers. In contrast, pMixFed introduces a novel, adaptive integration of layer-wise Mixup directly within the model parameter space—not the feature space. This integration is dynamically updated every communication round based on each client’s local accuracy (personalization) relative to the global accuracy (generalization). Governed by a tunable parameter $\mu$, pMixFed enables gradual, client-specific personalization across layers, mitigating global-local discrepancies and client drift in a fully dynamic manner. These innovations distinguish pMixFed from prior approaches, which rely on static strategies. We provided a theoretical analysis of pMixFed to study the properties of its convergence. Our experiments on three datasets demonstrated its superior performance over existing PFL methods. Empirically, pMixFed exhibited faster training times, increased robustness, and better handling of data heterogeneity compared to state-of-the-art PFL models. Future research directions include exploring multi-Modal personalization, exploiting other mixup variants within the parameter space and adapting pMixFed as a dynamic scheduling technique in resource-limited distributed setting.
Acknowledgments and Disclosure of Funding
This work was partly supported by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S. Department of Transportation National University Transportation Center) headquartered at Clemson University, Clemson, South Carolina, USA. Any opinions, findings, conclusions, and recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of TraCR, and the U.S. Government assumes no liability for the contents or use thereof.
@@ 8.1 Dynamic Mixup Degree
Among the various partitioning strategies for partial PFL discussed in the main paper and in pillutla2022federated, one of the most widely adopted techniques is assigning higher layers of local model $L_{l,k}^{t}$ to personalization while allowing the base layers $L_{g,k}^{t}$ to be shared across clients as the global model oh2021fedbabu; sun2023fedperfix; arivazhagan2019federated. This design aligns with insights from the Model-Agnostic Meta-Learning (MAML) algorithm collins2022maml, which demonstrates that lower layers generally retain task-agnostic, generalized features, while the higher layers capture task-specific, personalized characteristics. Accordingly, in this work, we designate the head of the model as containing personalized information, while the base layers represent generalized information shared across clients  [^6] It should be noted that our method is fully adaptable to different partial PFL designs as well discussed in pillutla2022federated..
8.1.1 Broadcasting:
To achieve a nuanced and gradual transition in the mixing process between the global model and the local model, we define the mix degree $\lambda_{i}$ as follows. The local model’s head $L_{n}$ remains frozen, and the head of the global model is excluded from being shared with the local model. The mix degree for each layer $\lambda_{i}$ increases incrementally based on the mix factor $\mu$, such that as we move toward the base layers, the personalization impact decreases. This dynamic behavior is represented as:
@@ 8.1.1 Broadcasting:
8.1.2 Aggregation:
The aggregation stage focuses on preserving the generalized information from the history of previous global models to mitigate the catastrophic forgetting problem. In contrast to the broadcasting process, the primary goal here is to retain the generalized information of the previous global model, which encapsulates the history of all prior models $H\big|_{1}^{t-1}{G}$. Therefore, the base layers should predominantly be shared from the global model, particularly during the first round of training, where the updated local models are still underdeveloped. Nonetheless, as we transition to the head, it’s less prominent to transfer knowledge from the previous global models. Figure 7 illustrates how the mixup degree transitions from the head to the base layers.
@@ 8.2 Adaptive Mix-Factor
To enable online, adaptive updates of the mix degree $\lambda_{i}$, we implemented several algorithms, as detailed in Section 5.4 Ablation Study. We observed a clear relationship between the mixup degree and the similar behavior to lr, which is further discussed in Section 4. Inspired by learning rate schedulers, we introduced the mix-factor $\mu$ to adaptively update the layer-wise mixup degree based on the current communication round $\delta=t/T$ and the relative performance of the current local model $L_{k}^{t}$ compared to the global model $G^{t}$. The Sigmoid function in Figure 7 illustrates how $\mu$ evolves with respect to $\delta$ and accuracy ($Acc$). The best results were achieved when setting $b=1$ and using the square of $Acc$ as an exponent. The rationale behind this approach is that a more experienced, better-performing model should share more information. Specifically, if the local model accuracy $Acc_{l}$ significantly exceeds that of the global model ($Acc_{G}$) in the current round ($t$) such that $Acc\gg 1$, less information is shared from the global model. Conversely, when $0<Acc<1$, the global model dominates the parameter updates in both the global and local models. Figure 8 shows the distribution of the calculated $\mu$ across different communication rounds for client 0 in the broadcasting stage. In the broadcasting stage, higher $Acc_{l}$ values result in freezing more layers for personalization, leading to a decrease in $\mu$. Conversely, during the aggregation phase, if the global model accuracy $Acc_{G}$ outperforms the updated local model accuracy $Acc_{l^{\prime}}^{(t+1)}$, more base layers are shared. This is especially important during the first round of training, where local models are less stable. As demonstrated in Section 6.3, our proposed method adapts dynamically to performance drops and high-distribution complexity, adjusting the mixup degree as needed. This adaptability is effective even when applied partially to partial PFL methods such as FedAlt, and FedSim.
@@ 8.3 Model Heterogeneity
8.4 Difference from other works:
While Mixup has previously been used in FL (e.g., XORMixup, FEDMIX, FedMix), its role was limited to data augmentation. Similarly, model partitioning has been explored in works like FedAlt and FedSim, but only with fixed cut-off layers. In contrast, pMixFed introduces a novel, adaptive integration of layer-wise Mixup directly within the model parameter space—not the feature space. This integration is dynamically updated every communication round based on each client’s local accuracy (personalization) relative to the global accuracy (generalization). Governed by a tunable parameter $\mu$, pMixFed enables gradual, client-specific personalization across layers, mitigating global-local discrepancies and client drift in a fully dynamic manner. These innovations distinguish pMixFed from prior approaches, which rely on static strategies.
@@ 8.4 Difference from other works:
9 Proof of Theorems:
Proof of Theorem 1 (SGD form with an effective step).
By definition of the pMixFed aggregation,
$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \sum_{k\in\mathcal{U}_{t}}\Omega_{k}\Big(\lambda_{k}^{t}\,\theta^{t}+(1-\lambda_{k}^{t})\,\theta_{k}^{t+1}\Big)\penalty 10000\ =\penalty 10000\ \bar{\lambda}^{\,t}\,\theta^{t}+(1-\bar{\lambda}^{\,t})\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\theta_{k}^{t+1},$
where $\bar{\lambda}^{\,t}:=\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\lambda_{k}^{t}\in[0,1]$. Each selected client performs one local SGD step from the broadcasted model $\theta^{t}$: $\theta_{k}^{t+1}=\theta^{t}-\eta_{\ell}\,g_{k}^{t}$, where $g_{k}^{t}$ is the stochastic gradient computed on client $k$ at $\theta^{t}$. Substituting into the aggregation gives
$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \bar{\lambda}^{\,t}\,\theta^{t}\penalty 10000\ +\penalty 10000\ (1-\bar{\lambda}^{\,t})\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\big(\theta^{t}-\eta_{\ell}g_{k}^{t}\big)\penalty 10000\ =\penalty 10000\ \theta^{t}-(1-\bar{\lambda}^{\,t})\,\eta_{\ell}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}.$
Therefore the server update is exactly an SGD step on $F(\theta)=\sum_{k}\Omega_{k}F_{k}(\theta)$ with the effective step size $\eta_{\mathrm{eff}}^{\,t}:=(1-\bar{\lambda}^{\,t})\,\eta_{\ell}$, i.e., $\theta^{t+1}=\theta^{t}-\eta_{\mathrm{eff}}^{\,t}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}$.
Proof of Theorem 2 (Coefficient matching with FedSGD).
Assume the server writes its update as a FedSGD step with server stepsize $\eta_{g}>0$: $\theta^{t+1}=\theta^{t}-\eta_{g}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}$. From Theorem 1 we also have the pMixFed SGD form $\theta^{t+1}=\theta^{t}-(1-\bar{\lambda}^{\,t})\,\eta_{\ell}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}$. Equality of the two updates for arbitrary realizations of the stochastic gradients holds if and only if their coefficients on the aggregated gradient coincide, i.e.,
[…]
Since by construction $\bar{\lambda}^{\,t}\in[0,1]$, we immediately obtain the admissible range $0\leq\eta_{g}/\eta_{\ell}\leq 1$. This proves the necessary and sufficient condition for coefficient matching.
Proof of Theorem 3 (Nonconvex rate).
Recall the pMixFed update in SGD form
$\theta^{t+1}\;=\;\theta^{t}\;-\;\eta_{\mathrm{eff}}\,\underbrace{\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}}_{:=\,\bar{g}^{t}},\qquad\eta_{\mathrm{eff}}\;\leq\;\frac{1}{L},$
where by (A2) the aggregated stochastic gradient $\bar{g}^{t}$ is unbiased, $\mathbb{E}[\bar{g}^{t}\mid\theta^{t}]=\nabla F(\theta^{t})$, and has bounded second moment $\mathbb{E}\|\bar{g}^{t}-\nabla F(\theta^{t})\|^{2}\leq\sigma^{2}$. By $L$-smoothness (A1),
$F(\theta^{t+1})\;\leq\;F(\theta^{t})+\big\langle\nabla F(\theta^{t}),\,\theta^{t+1}-\theta^{t}\big\rangle+\frac{L}{2}\,\|\theta^{t+1}-\theta^{t}\|^{2}.$
Substitute $\theta^{t+1}-\theta^{t}=-\eta_{\mathrm{eff}}\,\bar{g}^{t}$, take conditional expectation given $\theta^{t}$, and use $\mathbb{E}\|\bar{g}^{t}\|^{2}=\|\nabla F(\theta^{t})\|^{2}+\mathbb{E}\|\bar{g}^{t}-\nabla F(\theta^{t})\|^{2}\leq\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}$ to obtain
@@ 9 Proof of Theorems:
$\mathbb{E}\!\left[F(\theta^{t+1})\,\middle|\,\theta^{t}\right]\;\leq\;F(\theta^{t})\;-\;\eta_{\mathrm{eff}}\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\Big(\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}\Big).$
Rearranging,
$\mathbb{E}\!\left[F(\theta^{t+1})\,\middle|\,\theta^{t}\right]\;\leq\;F(\theta^{t})\;-\;\Big(\eta_{\mathrm{eff}}-\tfrac{L\eta_{\mathrm{eff}}^{2}}{2}\Big)\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
Since $\eta_{\mathrm{eff}}\leq 1/L$, we have $\eta_{\mathrm{eff}}-\tfrac{L\eta_{\mathrm{eff}}^{2}}{2}\geq\tfrac{\eta_{\mathrm{eff}}}{2}$; taking total expectation yields the one-step descent inequality
$\mathbb{E}\big[F(\theta^{t+1})\big]\;\leq\;\mathbb{E}\big[F(\theta^{t})\big]\;-\;\frac{\eta_{\mathrm{eff}}}{2}\,\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
Summing $(\star)$ over $t=0,\dots,T-1$ and telescoping gives
$\frac{\eta_{\mathrm{eff}}}{2}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\;\leq\;F(\theta^{0})-F^{\star}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,T\,\sigma^{2}.$
Divide both sides by $T\,\eta_{\mathrm{eff}}/2$ to conclude
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\;\leq\;\frac{2\,(F(\theta^{0})-F^{\star})}{T\,\eta_{\mathrm{eff}}}\;+\;L\,\eta_{\mathrm{eff}}\,\sigma^{2},$
which is precisely the claimed nonconvex rate.
Proof of Theorem 4 (Strongly convex case).
Assume (A1)–(A3) and $\eta_{\mathrm{eff}}\leq\min\{1/L,\,1/(2\mu)\}$. Starting again from the smoothness inequality and the SGD update,
$\mathbb{E}\big[F(\theta^{t+1})\,\big|\,\theta^{t}\big]\;\leq\;F(\theta^{t})\;-\;\eta_{\mathrm{eff}}\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\Big(\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}\Big).$
As before, with $\eta_{\mathrm{eff}}\leq 1/L$ we get
$\mathbb{E}\big[F(\theta^{t+1})\,\big|\,\theta^{t}\big]\;\leq\;F(\theta^{t})\;-\;\frac{\eta_{\mathrm{eff}}}{2}\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
Strong convexity implies the Polyak–Łojasiewicz (PL) inequality $\|\nabla F(\theta^{t})\|^{2}\geq 2\mu\big(F(\theta^{t})-F^{\star}\big)$. Plugging this into $(\dagger)$ and taking total expectation gives
$\mathbb{E}\big[F(\theta^{t+1})-F^{\star}\big]\;\leq\;\big(1-\mu\,\eta_{\mathrm{eff}}\big)\,\mathbb{E}\big[F(\theta^{t})-F^{\star}\big]\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
Unrolling the linear recursion,
$\mathbb{E}\big[F(\theta^{t})-F^{\star}\big]\;\leq\;(1-\mu\,\eta_{\mathrm{eff}})^{t}\big(F(\theta^{0})-F^{\star}\big)\;+\;\frac{L\,\eta_{\mathrm{eff}}\,\sigma^{2}}{2\mu}.$
Finally, strong convexity also yields $F(\theta)-F^{\star}\geq\tfrac{\mu}{2}\|\theta-\theta^{\star}\|^{2}$, so
$\mathbb{E}\,\|\theta^{t}-\theta^{\star}\|^{2}\;\leq\;\frac{2}{\mu}\,\mathbb{E}\big[F(\theta^{t})-F^{\star}\big]\;\leq\;(1-\mu\,\eta_{\mathrm{eff}})^{t}\,\|\theta^{0}-\theta^{\star}\|^{2}\;+\;\mathcal{O}\!\Big(\tfrac{\sigma^{2}}{\mu}\,\eta_{\mathrm{eff}}\Big),$
which is the claimed result.